# EvidenceMap: Unleashing the Power of Small Language Models with Evidence Analysis for Biomedical Question Answering

### v1

@@ 6 Related Work
6.1 Long-Form Question Answering
Long-Form Question Answering (LFQA) is a task in natural language processing that aims to generate detailed, coherent, and informative answers to address open-ended questions posed by users. Traditional research on LFQA focuses primarily on the method of generative models. ELI5 Fan et al. (2019) proposes an abstract generation model that integrates multi-task objective training. A following work Krishna et al. (2021) designs a system based on sparse attention and contrastive learning to address poor relevance issues. Recent studies concentrate more on retrieval-augmented long-form generation. One work Su et al. (2022) enhances the model with fine-grained key information relevant to the answers to improve the accuracy. WebCPM Qin et al. (2023) focuses on retrieval and synthesis, proposing an interactive web-based search that enables real-time interaction with search engines. Furthermore, evaluating the quality of LFQA is an important area. A study on LFQA evaluation Xu et al. (2023) employs a multidimensional assessment approach, focusing on aspects including factual, accuracy, and coherence. ASQA Stelmakh et al. (2022) proposes a reliable metric that can provide clear correctness standards to address the ambiguous definition of answer quality. Compared to previous studies, our research primarily focuses on the scenario of multiple and diverse evidence in biomedical domain, aiming to efficiently address LFQA tasks by enabling SLMs to explicitly learn evidence analysis.
6.2 Textual Knowledge Augmentation
Textual evidence, as an external knowledge, can significantly enhance the quality of language model generation in professional domains. Methods for knowledge augmentation can be categorized into two types. The first is the augmentation during inference. RAG Lewis et al. (2020), Self-RAG Asai et al. (2023), and FLARE Jiang et al. (2023) belong to this category, which improve generation effectiveness by retrieving relevant documents as factual knowledge and adjusting these texts during inference. Meanwhile, a model can reduce hallucinations from generation by incorporating reasoning knowledge, as demonstrated by CoT Wei et al. (2022), ToT Yao et al. (2024), and CoK Wang et al. (2023a), a series of sequential or structured reasoning steps are generated and executed explicitly. Furthermore, methods with an augmentation mixture of factual and reasoning knowledge can further combine the strengths of both, such as RAT Wang et al. (2024). Another type of augmentation is by tuning the model with domain text. Soft prompt tuning Lester et al. (2021); Yang et al. (2023) uses implicit learnable vectors to replace explicit tokens as prompts to flexibly adapt to multiple downstream tasks. Instruction tuning Wang et al. (2023b); Lin et al. (2023) improves models in knowledge-intensive tasks by constructing specific instruction datasets and learning the utilization of input information and contextual understanding. Compared to the aforementioned methods, our study proposes a novel knowledge augmentation process that prompts the inference of generative models by training the explicit evidence analysis abilities of a tiny pre-trained model.
@@ 7 Conclusion
In this work, we present a novel framework for generative biomedical question answering named EvidenceMap, which improves the ability of models to handle multiple and diverse evidence by explicitly training small language models in evidence analysis. The experimental results and cases show that EvidenceMap significantly improves the performance of the generative biomedical question answering by efficient utilization of evidence.
@@
First, the current evaluation of our method focuses solely on public datasets in the biomedical domain. Future evaluations will continue to examine its performance in other biomedical datasets, while also conducting tests in additional professional domains.
Second, this study has only tested a limited number of generative language models from the Llama 3 series, and further testing is needed on additional small generative models around the 3B scale, such as Phi-3.5-mini Abdin et al. (2024) and Qwen2.5-3B Yang et al. (2024).
Third, this work focuses primarily on the capabilities of SLMs while incorporating explicit evidence analysis. We will further explore the effects of learning evidence analysis on larger-scale pre-trained and generative models.
Fourth, the generalizability of learning evidence analysis skills and their potential transferability to other models remains an area for further exploration.
@@
Appendix A SLM and LLM Prompts
@@ A.1 Prompts for Evidence Map
The $\texttt{prompt}_{i}$ to obtain the representation of an individual evidence $X_{p}^{i}$ from an SLM:
@@
The $\texttt{prompt}_{e}$ to obtain the representation of the supportive evaluation of evidence $R_{eval}^{i}$ from an SLM:
@@
The $\texttt{prompt}_{c}$ to obtain the representation of the logical correlation of two pieces of evidence $R_{cor}^{(i,j)}$ from an SLM:
@@ Appendix C Details of Biomedical Datasets
The number of samples and the average amount of evidence per sample in the training and test sets from the two public biomedical datasets are shown in Table 5.
@@
Appendix D Results of Language Fluency
@@
Table 6: The comparison of language fluency measured by LLM on two datasets. Our SLM training architecture saves generation quality on fluency comparing with other language model based methods.

### v2

@@ 6 Related Work
6.1 Biomedical Evidence Analysis
Resolving issues in the biomedical domain requires a multifaceted analysis of related evidence. A recent study Singhal et al. (2025) suggests that addressing biomedical questions requires correct retrieval, accurate manipulation, and complete understanding of relevant knowledge. Some studies Jin et al. (2021); Singhal et al. (2023) indicate that models need to ignore irrelevant distractors to avoid incorrect interpretations of the evidence in the answers. Other studies focus on logical reasoning Chen et al. (2021) and summarization of evidence Xie et al. (2022); Zhang et al. (2024), thereby integrating medical evidence to support decision analysis and provide high-quality answers. Our study builds on these foundational works, integrating the above key aspects of evidence analysis, and propose a framework for evidence analysis and analysis-augmented generation to address biomedical question-answering tasks.
6.2 Textual Knowledge Augmented LLM
Textual evidence, as an external knowledge, can significantly enhance the quality of language model generation in professional domains. Methods for knowledge augmentation can be categorized into two types. The first is the augmentation during inference. RAG Lewis et al. (2020) and Self-RAG Asai et al. (2023) belong to this category, which improve generation effectiveness by retrieving relevant documents as factual knowledge and adjusting these texts during inference. Meanwhile, a model can reduce hallucinations from generation by incorporating reasoning knowledge, as demonstrated by CoT Wei et al. (2022) and ToT Yao et al. (2024), a series of sequential or structured reasoning steps are generated and executed explicitly. Furthermore, methods with an augmentation mixture of factual and reasoning knowledge can further combine the strengths of both, such as RAT Wang et al. (2024). Another type of augmentation is by tuning the model with domain text. Soft prompt tuning Lester et al. (2021); Yang et al. (2023) uses implicit learnable vectors to replace explicit tokens as prompts to flexibly adapt to multiple downstream tasks. Instruction tuning Wang et al. (2023b); Lin et al. (2023) improves models in knowledge-intensive tasks by constructing specific instruction datasets and learning the utilization of input information and contextual understanding. Compared to the aforementioned methods, our study proposes a novel knowledge augmentation process that prompts the inference of generative models by training the explicit evidence analysis abilities with a small pre-trained language model.
@@ 7 Conclusion
In this work, we present a novel framework for generative biomedical question answering named EvidenceMap, which efficiently addresses the robustness issues of language model generation with multiple pieces of evidence. Inspired by evidence analysis in the biomedical domain, our framework explicitly trains a small language model on multifaceted aspects of evidence to prompt the generation of answers. The experimental results and representative cases on public datasets show that EvidenceMap can significantly surpass the RAG method with equivalent and larger language models by fine-tuning a 66M pre-trained model to prompt a generative model around 3B.
@@
First, we currently evaluate our method with popular long-form metrics on public biomedical datasets. Future studies continue to propose specific and rational metrics for the biomedical domain, while also conducting tests in additional professional domains.
Second, the evidence map and the learning of evidence analysis that we propose provide a guide to improve the ability to use evidence. We will continue to explore enabling the model to analyze fine-grained and symbolic logical relationships among evidence to further improve model robustness while using multiple pieces of evidence.
Third, despite our method demonstrating superior performance in small language models, we will continue to conduct in-depth research on the interpretability and visualization of model inference, exploring further advantages in evidence analysis.
Fourth, the generalizability of learning evidence analysis and the potential transferability to other models still remains an area for further exploration.
@@
Appendix A Prompts Used in EvidenceMap
@@ A.1 Prompts for Evidence Map
The $\texttt{prompt}_{i}$ to obtain the representation of an individual evidence $X_{p}^{i}$ from an small language model:
@@
The $\texttt{prompt}_{e}$ to obtain the representation of the supportive evaluation of evidence $R_{eval}^{i}$ from an small language model:
@@
The $\texttt{prompt}_{c}$ to obtain the representation of the logical correlation of two pieces of evidence $R_{cor}^{(i,j)}$ from an small language model:
@@ Appendix C Details of Biomedical Datasets
The number of samples and the average amount of evidence per sample in the training and test sets from the two public biomedical datasets are shown in Table 5. The obvious difference in the number of evidence and the results shown in Table 1 indicates that our method can be fitted with variable data conditions.
@@
Appendix D Results on More Generative Models
We also evaluate our method on more small generative models around the 3B parameter size, including Qwen-2.5-3B-Instruct Yang et al. (2024) and Phi-3.5-Mini-Instruct Abdin et al. (2024). The results of experiments on BioASQ dataset shown in Figure 6 demonstrate that the performance of EvidenceMap can remain consistent across a wider range of small generative models.
Figure 6: Results on BioASQ with various small generative language models which have parameter size around 3B. The best scores of three metrics are highlighted.
Appendix E Results of Language Fluency
@@
Table 6: The comparison of language fluency measured by LLM on two datasets. Our language model training architecture saves generation quality on fluency comparing with other language model based methods.